# Bounds on the number of squares in recurrence sequences

### v2

@@ 4.6. Small $d$
If $-N_{\alpha}$ is an even square, $1\leq u\leq 4$ and $y_{\pm 1}<\max\left(1,\frac{76\left|N_{\alpha}\right|^{3/2}}{\sqrt{d}\left(|g|\mathcal{N}_{d^{\prime},4}\right)^{2}}\right)$, then we must have $u=1$, $t^{2}-du^{2}=-4$, $N_{\alpha}\equiv 12\pmod{16}$ and $\gcd\left(a^{2},d\right)=1,4$.
@@
(1-i) We show that $u=4$, $-N_{\alpha}$ an even square and $t^{2}-du^{2}=4$ is not possible.
@@
(1-ii) We show that $u=3$, $-N_{\alpha}$ an even square and $t^{2}-du^{2}=4$ is not possible.
@@
(1-iii) We show that $u=2$, $-N_{\alpha}$ an even square and $t^{2}-du^{2}=4$ is not possible.
@@
(1-iv) We show that $u=1$, $-N_{\alpha}$ an even square and $t^{2}-du^{2}=4$ is not possible.
@@
(2-i) We show that if $u=1$, $-N_{\alpha}$ is an even square and $t^{2}-du^{2}=-4$, then $N_{\alpha}\equiv 12\pmod{16}$.
We find that $y_{\pm 1}\equiv 0,1,4,9\pmod{16}$ is not possible if $-N_{\alpha}=b_{1}^{2}$ where $b_{1}\equiv 4\pmod{8}$.
@@
For $d<80$ with $u=1$ and $-N_{a}$ an even square divisble by $64$, there is just one possibility: $d=68$, $t=8$, $u=1$, $a=2$, so $N_{\alpha}=-64$ and $y_{-1}=25$. But here, $\frac{76\left|N_{\alpha}\right|^{3/2}}{\left(\sqrt{d}\left(|g|\mathcal{N}_{d^{\prime},4}\right)^{2}\right)}<\left(76/16^{2}\right)\left|N_{\alpha}\right|^{3/2}/\sqrt{d}<19$, so this case is excluded too and hence we can exclude $8|b_{1}$ altogether.
Hence $b_{1}\equiv 2\pmod{4}$ and so $-N_{\alpha}\equiv 4\pmod{16}$.
(2-ii) We show that if $u=1$, $-N_{\alpha}$ is an even square, $t^{2}-du^{2}=-4$, and $N_{\alpha}\equiv 12\pmod{16}$, then $\gcd\left(a^{2},d\right)=1$ or $4$.
We have $v_{2}\left(\gcd\left(a^{2},d\right)\right)=0$ or $2$. Since $-N_{\alpha}\equiv 4\pmod{16}$, we have $v_{2}\left(b^{\prime}\right)=2$ in the first case and 00 in the second case.
We consider first $\gcd\left(a^{2},d\right)>4$ odd. So by Lemma 3.12 and the argument in the previous paragraph, $|g|\mathcal{N}_{d^{\prime},4}=4\sqrt{\gcd\left(a^{2},d\right)}$. Hence $\frac{76\left|N_{\alpha}\right|^{3/2}}{\sqrt{d}\left(|g|\mathcal{N}_{d^{\prime},4}\right)^{2}}=\frac{76\left|N_{\alpha}\right|^{3/2}}{16\sqrt{d}\gcd\left(a^{2},d\right)}$. If $\gcd\left(a^{2},d\right)\geq 25$, then the right-hand side is less than $40.19\left|N_{\alpha}\right|<\left|N_{\alpha}\right|/$ 4, the right-hand side being the lower bound for $y_{\pm 1}$ from Lemma 3.5(c).
@@
Now consider $\gcd\left(a^{2},d\right)>4$ even. Then $v_{2}\left(\gcd\left(a^{2},d\right)\right)=2$ (so $\gcd\left(a^{2},d\right)\equiv 4\pmod{16}$). So by Lemma 3.12 and the argument above, $|g|\mathcal{N}_{d^{\prime},4}=2\sqrt{\gcd\left(a^{2},d\right)}$. Hence $\frac{76\left|N_{\alpha}\right|^{3/2}}{\sqrt{d}\left(|g|\mathcal{N}_{d^{\prime},4}\right)^{2}}=\frac{76\left|N_{\alpha}\right|^{3/2}}{4\sqrt{d}\gcd\left(a^{2},d\right)}$.
@@
(2-iii) We show that $u=2$, $-N_{\alpha}$ an even square and $t^{2}-du^{2}=-4$ is not possible.
@@
Suppose that $d\geq 2$. Then $t^{2}=4d-4$, so $t^{2}\geq 2d$. If $a^{2}<0.292d$, then $t-a>\sqrt{2d}-\sqrt{0.292d}$ and so $(t-a)^{2}/4>0.19d>0.19\left|N_{\alpha}\right|$. We can write $4y_{\pm 1}=\left(t\pm au\right)^{2}-N_{\alpha}u^{2}$. So here with $u=2$, we have $y_{\pm 1}\geq\left(t-a\right)^{2}/4+\left|N_{\alpha}\right|>1.19\left|N_{\alpha}\right|$. But $76\left|N_{\alpha}\right|/64=1.1875\left|N_{\alpha}\right|$. So we can exclude the case where $u=2$, $-N_{\alpha}$ an even square and $t^{2}-du^{2}=-4$. ∎
@@ 1 $41$ [0, -1, -9, 11] [41, 71, 80753, 470861]
Let $n\geq 5$ be an odd integer. Put $a=\left(n^{2}-9\right)/4$, $d=\left(n^{4}-2n^{2}+17\right)/16$, so that $N_{\alpha}=4-n^{2}$. With $t=\left(n^{2}-1\right)/2$ and $u=2$, $\varepsilon=\left(t+u\sqrt{d}\right)/2$ is a unit in $\mathcal{O}_{\mathbb{Q}\left(\sqrt{d}\right)}$. We have $y_{1}=\left(n^{2}-3\right)^{2}/4$ and $y_{-1}=n^{2}$. This example shows that Conjecture 1.2 is best possible, if true.
@@
From (3.14) with $b=u=1$, we find that $(2t+a)^{2}-8y_{1}=a^{2}-8$. Since $-N_{\alpha}=d-a^{2}=t^{2}+4-a^{2}$ is a square, we put $a=2$, so if $y_{1}$ is a square, then $22t+$ 2 is an element of the sequence $\left(t_{n}\right)_{n\geq 0}$ with $t_{0}=6$, $t_{1}=238$ and $t_{n+1}=34u_{n}-u_{n-1}$. The corresponding values of $\sqrt{y_{1}}$ are the elements of the sequence $\left(u_{n}\right)_{n\geq 0}$ with $u_{0}=5$, $u_{1}=169$ and $u_{n+1}=34u_{n}-u_{n-1}$. Then $y_{1}$ is approximately $t^{2}/2\approx\left|N_{\alpha}\right|/2\approx\left|N_{\alpha}\right|^{3/2}/\left(2\sqrt{d}\right)$.
@@
We can show that the lower bound in Lemma 3.5(c) is actually best possible. Let $n$ be a positive integer such that $a=2n^{2}-3$ is not divisible by $5$ and put $d=4n^{4}-8n^{2}+8$. We have $N_{\alpha}=1-4n^{2}$. Then $(t,u)=(a+1,1)$, $y_{-1}=n^{2}$ and so $y_{-1}/\left|N_{\alpha}\right|\rightarrow 1/4$ from above as $n\rightarrow+\infty$. There exist families with the same limit for $y_{-1}/\left|N_{\alpha}\right|$ when $-N_{\alpha}$ is a square too.
The sequences given in Lemma 3.6 show that for $k=-1$, the lower bound in Lemma 3.5(c) can be attained too. At least, when the lower bound is $1$.

### v3

@@ 4.6. Small $d$
If $N_{\alpha}$ is an even square, $1\leq u\leq 4$ and $y_{\pm 1}<\max\left(1,\frac{76\left|N_{\alpha}\right|^{3/2}}{\sqrt{d}\left(|g|\mathcal{N}_{d^{\prime},4}\right)^{2}}\right)$, then we must have $u=1$, $t^{2}-du^{2}=-4$, $N_{\alpha}\equiv 12\pmod{16}$ and $\gcd\left(a^{2},d\right)=1,4$.
@@
(1-i) We show that $u=4$, $N_{\alpha}$ an even square and $t^{2}-du^{2}=4$ is not possible.
@@
(1-ii) We show that $u=3$, $N_{\alpha}$ an even square and $t^{2}-du^{2}=4$ is not possible.
@@
(1-iii) We show that $u=2$, $N_{\alpha}$ an even square and $t^{2}-du^{2}=4$ is not possible.
@@
(1-iv) We show that $u=1$, $N_{\alpha}$ an even square and $t^{2}-du^{2}=4$ is not possible.
@@
(2-i) We show that if $u=1$, $N_{\alpha}$ is an even square and $t^{2}-du^{2}=-4$, then $N_{\alpha}\equiv 12\pmod{16}$.
We find that $y_{\pm 1}\equiv 0,1,4,9\pmod{16}$ is not possible if $N_{\alpha}=b_{1}^{2}$ where $b_{1}\equiv 4\pmod{8}$.
@@
For $d<80$ with $u=1$ and $N_{a}$ an even square divisble by $64$, there is just one possibility: $d=68$, $t=8$, $u=1$, $a=2$, so $N_{\alpha}=-64$ and $y_{-1}=25$. But here, $\frac{76\left|N_{\alpha}\right|^{3/2}}{\left(\sqrt{d}\left(|g|\mathcal{N}_{d^{\prime},4}\right)^{2}\right)}<\left(76/16^{2}\right)\left|N_{\alpha}\right|^{3/2}/\sqrt{d}<19$, so this case is excluded too and hence we can exclude $8|b_{1}$ altogether.
Hence $b_{1}\equiv 2\pmod{4}$ and so $N_{\alpha}\equiv 4\pmod{16}$.
(2-ii) We show that if $u=1$, $N_{\alpha}$ is an even square, $t^{2}-du^{2}=-4$, and $N_{\alpha}\equiv 12\pmod{16}$, then $\gcd\left(a^{2},d\right)=1$ or $4$.
We have $v_{2}\left(\gcd\left(a^{2},d\right)\right)=0$ or $2$. Since $N_{\alpha}\equiv 4\pmod{16}$, we have $v_{2}\left(b^{\prime}\right)=2$ in the first case and 0 in the second case.
We consider first $\gcd\left(a^{2},d\right)>4$ odd. So by Lemma 3.12 and the argument in the previous paragraph, $|g|\mathcal{N}_{d^{\prime},4}=4\sqrt{\gcd\left(a^{2},d\right)}$. Hence $\frac{76\left|N_{\alpha}\right|^{3/2}}{\sqrt{d}\left(|g|\mathcal{N}_{d^{\prime},4}\right)^{2}}=\frac{76\left|N_{\alpha}\right|^{3/2}}{16\sqrt{d}\gcd\left(a^{2},d\right)}$. If $\gcd\left(a^{2},d\right)\geq 25$, then the right-hand side is less than $0.19\left|N_{\alpha}\right|<\left|N_{\alpha}\right|/4$, the right-hand side being the lower bound for $y_{\pm 1}$ from Lemma 3.5(c).
@@
[…]
@@
(2-iii) We show that $u=2$, $N_{\alpha}$ an even square and $t^{2}-du^{2}=-4$ is not possible.
@@
Suppose that $d\geq 2$. Then $t^{2}=4d-4$, so $t^{2}\geq 2d$. If $a^{2}<0.292d$, then $t-a>\sqrt{2d}-\sqrt{0.292d}$ and so $(t-a)^{2}/4>0.19d>0.19\left|N_{\alpha}\right|$. We can write $4y_{\pm 1}=\left(t\pm au\right)^{2}-N_{\alpha}u^{2}$. So here with $u=2$, we have $y_{\pm 1}\geq\left(t-a\right)^{2}/4+\left|N_{\alpha}\right|>1.19\left|N_{\alpha}\right|$. But $76\left|N_{\alpha}\right|/64=1.1875\left|N_{\alpha}\right|$. So we can exclude the case where $u=2$, $N_{\alpha}$ an even square and $t^{2}-du^{2}=-4$. ∎
@@ 1 $41$ [0, -1, -9, 11] [41, 71, 80753, 470861]
Let $n\geq 5$ be an odd integer. Put $a=\left(n^{2}-9\right)/4$ and $d=\left(n^{4}-2n^{2}+17\right)/16$, so that $N_{\alpha}=4-n^{2}$. With $t=\left(n^{2}-1\right)/2$ and $u=2$, $\varepsilon=\left(t+u\sqrt{d}\right)/2$ is a unit in $\mathcal{O}_{\mathbb{Q}\left(\sqrt{d}\right)}$. We have $y_{1}=\left(n^{2}-3\right)^{2}/4$ and $y_{-1}=n^{2}$. This example shows that Conjecture 1.2 is best possible, if true.
@@
From (3.14) with $b=u=1$, we find that $(2t+a)^{2}-8y_{1}=a^{2}-8$. Since $N_{\alpha}=d-a^{2}=t^{2}+4-a^{2}$ is a square, we put $a=2$, so if $y_{1}$ is a square, then $2t+2$ is an element of the sequence $\left(t_{n}\right)_{n\geq 0}$ with $t_{0}=6$, $t_{1}=238$ and $t_{n+1}=34u_{n}-u_{n-1}$. The corresponding values of $\sqrt{y_{1}}$ are the elements of the sequence $\left(u_{n}\right)_{n\geq 0}$ with $u_{0}=5$, $u_{1}=169$ and $u_{n+1}=34u_{n}-u_{n-1}$. Then $y_{1}$ is approximately $t^{2}/2\approx\left|N_{\alpha}\right|/2\approx\left|N_{\alpha}\right|^{3/2}/\left(2\sqrt{d}\right)$.
@@
We can show that the lower bound in Lemma 3.5(c) is actually best possible. Let $n$ be a positive integer such that $a=2n^{2}-3$ is not divisible by $5$ and put $d=4n^{4}-8n^{2}+8$. We have $N_{\alpha}=1-4n^{2}$. Then $(t,u)=(a+1,1)$, $y_{-1}=n^{2}$ and so $y_{-1}/\left|N_{\alpha}\right|\rightarrow 1/4$ from above as $n\rightarrow+\infty$. There exist families with the same limit for $y_{-1}/\left|N_{\alpha}\right|$ when $N_{\alpha}$ is a square too.
The sequences given in Lemma 3.6 show that for k=1k=-1, the lower bound in Lemma 3.5(c) can be attained too. At least, when the lower bound is $1$.